bernard:

Click an item in the list or group of pictures at the bottom of the problem and, holding the button down, drag it into the correct position in the answer box. Release your mouse button when the item is place. If you change your mind, drag the item to the trashcan. Click the trashcan to clear all your answers. Enter the equation of the circle described below. Center (3, -2), radius = 5

IsabelFriedmannbells:

@bernard Where are the options at?

XioGonz:

Do you have a picture to accompany this post?

bernard:

1 attachment
bernard:

1 attachment
XioGonz:

Thank you

IsabelFriedmannbells:

@bernard what is this math about??? im not understanding how i need to help you

XioGonz:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @IsabelFriedmannbells @bernard what is this math about??? im not understanding how i need to help you \(\color{#0cbb34}{\text{End of Quote}}\) If you can't help him, leave. No need to reply on a post that you have no idea how to help in.

bernard:

It's geometry @isabelFriedmanbells this is Geometry

XioGonz:

Alright, well you have to use the SF of the equation of a circle

XioGonz:

DO you know what the general equation is?

bernard:

I dont it have one TBH they jus gave me this out the blue tbh because of corona we not being taught nothing they just gave us compters and sent us home before Corona got bad bad

XioGonz:

Okay well the equation is \[(x-h)^2+(y-k)^2=r^2\]

bernard:

THank you

XioGonz:

That's not the answer though

XioGonz:

Center=(h,k) which equals to (3,-2)

XioGonz:

I used r for the radius and r=5

XioGonz:

\[(x - 3)^2 + (y - (-2))^2 = (5)^2\]

XioGonz:

right now what we're doing is placing the values into the general equation

XioGonz:

\[(x -3)^2 + (y + 2)^2 = 25\]

XioGonz:

\[(x^2 - 6. + 9) + (y^2 + 4y + 4) = 25\]

Serenity1jacksonsl:

i thaught you could not give direct answers

XioGonz:

\[x^2 +y^2 - 6x + 4y +9+4 -25= 0\]

XioGonz:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Serenity1jacksonsl i thaught you could not give direct answers \(\color{#0cbb34}{\text{End of Quote}}\) I am not giving a direct answer

XioGonz:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @XioGonz \[x^2 +y^2 - 6x + 4y +9+4 -25= 0\] \(\color{#0cbb34}{\text{End of Quote}}\) I'll let you do the last step on your own

bernard:

Okay im still working on it thank you

XioGonz:

No problem